# On Mutation-Guided Unit Test Generation

### v2

@@ Abstract
Keywords Unit Test Generation  $\cdot$ Mutation Testing  $\cdot$ Large Language Models
1 Introduction
@@
Unit test generation is a fundamental task in software engineering, crucial for ensuring code quality and reducing manual testing effort (Naik and Tripathy, 2011; Fraser and Arcuri, 2014; Corno et al., 2004; Anand et al., 2013). With the rapid advancement of Large Language Models (LLMs), research has shifted from traditional techniques—such as random testing (Pacheco and Ernst, 2007; Hamlet, 1994; Andrews et al., 2011; Oriat, 2005; Pacheco et al., 2007) and search-based methods (Fraser and Arcuri, 2011; Rojas et al., 2016, 2015; Vivanti et al., 2013)—to LLM-based approaches for generating unit tests (Fan et al., 2023; Wang et al., 2024; Hou et al., 2024).
@@
However, most existing work evaluates generated test cases primarily based on code coverage metrics, such as line or branch coverage. According to prior studies (Chekam et al., 2017; Andrews et al., 2006; Inozemtseva and Holmes, 2014), high code coverage does not necessarily imply strong fault detection capability. Other research (Chekam et al., 2017; Alshahwan et al., 2024; Foster et al., 2025) suggests that mutation score is a more reliable and meaningful metric for evaluating the effectiveness of test cases. Mutation testing introduces artificial faults, known as mutants, into the program by applying syntactic changes using mutation operators. The generated test cases are executed against the mutants, and if a test causes a mutant to fail (i.e., “kills” it), it is considered effective. The mutation score is the ratio of killed mutants over the total number of generated mutants.
@@
Recent efforts have begun exploring mutation-based evaluation in the context of LLM-generated test cases. For example, Foster et al. (Foster et al., 2025) investigate regression testing scenarios in privacy-critical Kotlin applications such as WhatsApp and Instagram. They focus on generating test cases using LLMs based on existing ones, and on generating mutants using LLMs, which introduces challenges such as detecting equivalent mutants that no test can kill. Another work by Dakhel et al. (Dakhel et al., 2024) targets Python benchmarks and proposes a prompt-based approach to improve mutation score. However, they iteratively add a single mutant until they cannot kill the new mutant. Therefore, they do not attempt to converge towards a maximum mutation score to maximize fault detection. Moreover, their study does not analyze the reasons why live mutants remain undetected, nor does it consider uncovered mutants, which are challenges that we address in our study.
@@
While these efforts represent promising first steps, the use of mutation testing as a core evaluation and improvement strategy for LLM-generated test cases remains underexplored. To address this gap, we propose MutGen, a mutation-guided LLM-based approach that explicitly aims to maximize mutation score. Specifically, a key feature of MutGen is that we incorporate mutation feedback (i.e., live and uncovered mutant information) into prompt construction to guide the LLM in generating more effective test cases. Further, as discussed in prior studies (Tian et al., 2025; Wang et al., 2025a; Pan et al., 2024), incorporating a fixing step is crucial for improving test generation effectiveness. While prior work primarily focuses on improving metrics such as code coverage or test passing rate, our work explicitly targets the fault detection capability of test cases, measured by the mutation score. To this end, our approach introduces a fixing step that specifically addresses assertion failures in the generated test cases.
@@
Moreover, we observe that comments in the target code can mislead LLMs into producing irrelevant or incorrect tests. To address this, we introduce a code summarization step during preprocessing that replaces misleading comments with concise summaries, which are then embedded into the prompt. Last, since achieving maximum mutation score is inherently difficult, we further introduce an iterative test generation strategy to push LLMs toward killing more live mutants. Once the mutation score converges, we analyze both uncovered and live mutants, as well as specific mutation operators that remain particularly challenging for the LLM.
@@
To evaluate MutGen, we use two datasets: HumanEval-Java, widely adopted in prior work (Tian et al., 2025; Dakhel et al., 2024; Siddiq et al., 2024), and a new LeetCode-Java corpus containing 100 randomly selected algorithmic problems with ground-truth solutions. The experimental results show that test cases generated by MutGen outperform both EvoSuite, a state-of-the-art search-based technique, and a vanilla LLM-based method in terms of mutation score. MutGen achieves mutation scores of 89.5% and 89.1% on HumanEval-Java and LeetCode-Java, respectively, across a total of 1,144 and 1,900 mutants.
@@
2 Motivating Example and Informal Description
@@
When combined with the example code (including comments), LLMs can generate test cases that achieve high line and branch coverage. However, as demonstrated in prior work (Foster et al., 2025; Chekam et al., 2017; Alshahwan et al., 2024), high coverage does not necessarily imply strong fault detection capability, for example when measured as mutation score. For instance, in our experiments, LLMs generate tests for the subject id_81 from HumanEval-Java with 100% line and branch coverage, yet the corresponding mutation score is only 4%.
Listing 2 shows a representative test function generated by Llama-3.3. In practice, multiple such test functions may be generated, often with a variety of date inputs aimed at maximizing code coverage. However, we observe that most of these inputs fail to cover corner cases such as February 29th on leap years, likely due to the absence of such examples in the provided comment block. The comments in the target code may also introduce additional issues. For example, the comment begins with “You have to write a function…”, which, when included as part of the prompt, may mislead the LLM into replicating the target code logic rather than generating a meaningful test function. Additionally, Llama-3.3 occasionally produces dates in incorrect formats such as dd-mm-yyyy, likely due to ambiguity or insufficient specification of the expected format in the comments provided within the program under test.
@@
which fails in both for the original function and mutant. However, it frequently fails to generate:
@@
1@Test
2public void testDayBoundaryInApril() {
3 String date = "04-01-2025";
4 assertTrue(ValidDate.validDate(date));
5}
@@ 3 String date = "04-01-2025";
To address these issues, we propose MutGen, an approach that improves the fault detection capability of LLM-generated test cases via a two-stage process: preprocessing and generation, which are detailed in Section 3.2 and Sections 3.3-3.5, respectively.
2.1 Preprocessing Stage
@@
“This Java function, ‘validDate‘, validates a given date string in the format "mm-dd-yyyy" and returns ‘true‘ if the date is valid according to specific rules regarding month and day ranges, and ‘false‘ otherwise. The input date string must be non-empty and follow the exact specified format. If any of these conditions are not met, the function will return ‘false‘.”
To address the second issue, we leverage mutation feedback to guide the LLM in generating test cases with higher fault detection capability. Before the generation stage, MutGen extracts mutation feedback from mutation reports produced by mutation testing tools. In this example, we use PITest (pit, Accessed: 2025) to generate a mutation report, as shown in Figure 3. From the report, MutGen retrieves key information, including the mutation location (line number), the mutant status (killed, survived, or uncovered), and the mutation operator applied to create the mutant. MutGen then integrates this feedback with the source code, forming a prompt (shown in Figure 4) that guides the LLM in generating more effective test cases during the generation stage. For example, the mutation feedback extracted from the report for the mutant shown in Listing 1 is as follows:
@@
2.2 Generation Stage
Prompt Used for Generation # Unit Test Generator for Java using Junit4 ## Source Code Summary {{summary}} ## Source File {{source code without comments}} ## Mutation Feedback Analyze the live mutant to think how to find the defects, the generated test cases must look very different with the existing ones. 21. conditional boundary at line 24, uncovered: if (day<=1||day>30) return false; ## Output Format in YAML new_tests: - test_behavior: … test_name: … test_code: … new_imports: …
@@ 2.2 Generation Stage
To further enhance effectiveness, MutGen introduces an iterative generation strategy. In each iteration, additional test cases are generated to target the remaining live and uncovered mutants that were not killed by the previous test suite. After two iterations, for the running example, MutGen achieves a 100% mutation score. During this process, LLMs may occasionally generate test cases with method names that conflict with those already created. For example, as shown in Listing 2, the LLM generates a test case for the boundary value 04-01-2025 in April. Later, it may generate another test case for 04-31-2025 also intended to test boundary conditions. However, both may be assigned the same method name, leading to naming conflicts. To resolve this, MutGen incorporates an additional guideline, specifically addressing method name conflicts, into the prompt used for fixing (shown in Figure 5).
In contrast, as demonstrated in Section 4, for this subject in HumanEval-Java, using the standard prompt introduced at the beginning of this section, the LLM can only produce a test suite achieving a 53% mutation score, which remains unchanged even after four iterations.
Prompt Used for Fixing ## Failed Test The following test failed in the test suite: {{ failed_test with error message }} You are processing execution failures, please match only one of these guidelines and then try to correct the failed tests according to the given error messages. 1. Direct Invocation: • Always use assertion functions directly (e.g., assertEquals(value1, value2)) instead of prefixed calls like Assert.assertEquals(value1, value2). 2. Handling assertEquals Mismatches: • If an error follows the pattern: • test_file_name:line_number: expected:error_value but was:correct_value • And involves assertEquals, replace the expected value with the correct one. • Example: assertEquals(error_value, xxx); → assertEquals(correct_value, xxx); 3. Fixing assertTrue / assertFalse Errors: • If an error follows the pattern: • classname:line_number:null • And involves assertTrue or assertFalse, swap the function name while keeping all parameters unchanged. • Example: assertTrue(xxx); → assertFalse(xxx); 4. Handling ambiguous • If an error has the pattern: • reference to assertEquals is ambiguous • And involves assertEquals, add type casts to both parameters based on their respective types. • Example: assertEquals(expectedValue, actualValue); → assertEquals((long) expectedValue, (long) actualValue); 5. Method name conflicts • If an error follows the pattern: method is already defined in the class, • only append a number to the method name. • Example: public void test_null() → public void test_null1(); 6. Other errors • Try to fix the errors according to the error messages. ## Output Format in YAML The same applied to Figure 4.
@@
3 Our Approach
In this section, we precisely describe and formalize our proposed approach, MutGen. As shown in Figure 6, MutGen takes a source program as input and returns test cases aimed at improving its fault-detection effectiveness by maximizing the test suite’s mutation score. We highlight all data flows in red in the figure. Steps 1, 2, and 5 are described in Sections 3.2, 3.3, and 3.4, respectively. We do not include Section 3.5 as it reuses the components already shown in the figure. We begin with a high-level overview of MutGen, followed by detailed and formal descriptions of each individual component.
3.1 Overview
Let =$\mathcal{M}=\{m_{1},m_{2},\dots,m_{n}\}$ denote the set of mutants generated from the program under test. The mutation score $\mu(T)$ of a test suite $T$ is defined as:
@@ 3.1 Overview
The preprocessing stage prepares information to guide test case generation, including code summarization and mutation feedback, as detailed in Section 3.2. In the generation stage, the LLM is prompted with mutation feedback, such as descriptions of mutation operators and diff-like representations of mutants, to generate test cases.
Figure 6: MutGen Overview
Let =$\mathcal{R}=\{r_{1},r_{2},\dots,r_{k}\}$ denote the set of mutation feedback entries (i.e., killed, live, uncovered), where each $r_{i}$ is a tuple $\langle\text{id}_{i},m_{i},s_{i},\text{op}_{i}\rangle$, respectively representing the mutant identifier, the mutated statement, the mutant status, and the applied mutation operator. MutGen first generates an initial test suite $\mathcal{T}$:
@@
After executing $\mathcal{T}$, MutGen collects a set =$\mathcal{F}=\{f_{1},...,f_{q}\}$ consisting of all test cases that encountered execution failures and corresponding error messages. For each $f_{i}$, MutGen constructs a second prompt $\text{Prompt}_{\text{fail}}$ (as illustrated in the example shown in Figure 5) and invokes the LLM to generate a set of fixed test cases $\mathcal{T}_{fix}$:
@@
This process will be detailed in Section 3.4.
@@
Our assumption, as reported in the mutation testing literature (Chekam et al., 2017; Andrews et al., 2006), is that by maximizing the mutation score we also maximize fault detection effectiveness. In any case, achieving mutation coverage is known to be a more stringent criterion that line or branch coverage (Foster et al., 2025) and therefore enable the generation of more effective test suites. However, as revealed in recent work (Gao et al., 2025), finding an optimal prompt for LLMs remains a significant challenge and we introduce an approximate and iterative solution, detailed in Section 3.5, which we investigate empirically and whose limitations we analyze, for example in terms of mutation operators that are difficult to kill.
3.2 Preprocessing
@@
Code Summarization. As illustrated in Figure 6 and the example in Section 2.1, the prompt instructs the LLM to summarize the purpose and the input format of the target code, resulting in a natural language description denoted as $desc$.
@@
As shown in the example in Section 2.1, the corresponding mutation feedback entry is represented as:
@@
3.3 Prompting with Mutation Feedback

### v4

# Mutation-Guided Unit Test Generation with a Large Language Model

@@ Abstract
Unit Test Generation, Mutation Testing, Large Language Models
I Introduction
@@
Unit test generation is a fundamental task in software engineering, crucial for ensuring code quality and reducing manual testing effort [48, 30, 22, 10]. With the rapid advancement of Large Language Models (LLMs), research has shifted from traditional techniques—such as random testing [53, 36, 12, 50, 54] and search-based methods [29, 59, 58, 65]—to LLM-based approaches for generating unit tests [27, 66, 37].
@@
However, most existing work evaluates generated test cases primarily based on code coverage metrics, such as line or branch coverage. According to prior studies [18, 11, 39], high code coverage does not necessarily imply strong fault detection capability. Other research [18, 9, 28] suggests that the mutation score is a more reliable and meaningful metric for evaluating the effectiveness of test cases. Mutation testing introduces artificial faults, known as mutants, into the program by applying syntactic changes using mutation operators. The generated test cases are executed against the mutants, and if a test causes a mutant to fail (i.e., “kills” it), it is considered effective. The mutation score is the ratio of killed mutants to the total number of generated mutants.
@@
Recent efforts have begun exploring mutation-based evaluation in the context of LLM-generated test cases. For example, Foster et al. [28] investigate regression testing scenarios in privacy-critical Kotlin applications such as WhatsApp and Instagram. They focus on generating test cases using LLMs based on existing ones, and on generating mutants using LLMs, which introduces challenges such as detecting equivalent mutants that no test can kill. Another work by Dakhel et al. [23] targets Python benchmarks and proposes a prompt-based approach to improve mutation score. However, they iteratively add a single mutant until they cannot kill the new mutant. Therefore, they do not attempt to converge towards a maximum mutation score to maximize fault detection. Moreover, their study does not analyze the reasons why live mutants remain undetected, nor does it consider uncovered mutants, which are challenges that we address in our study.
@@
While these efforts represent promising first steps, the use of mutation testing as a core evaluation and improvement strategy for LLM-generated test cases remains underexplored. To address this gap, we propose MutGen, a mutation-guided LLM-based approach that explicitly aims to maximize mutation score. Specifically, a key feature of MutGen is that we incorporate mutation feedback (i.e., live and uncovered mutant information) into prompt construction to guide the LLM in generating more effective test cases. Further, as discussed in prior studies [63, 68, 55], incorporating a fixing step is crucial for improving test generation effectiveness. While prior work primarily focuses on improving metrics such as code coverage or test passing rate, our work explicitly targets the fault detection capability of test cases, measured by the mutation score. To this end, our approach introduces a fixing step that specifically addresses assertion failures in the generated test cases, caused by incorrect usage of assertion functions, including erroneous invocations and the use of inappropriate parameters.
@@
Moreover, we observe that comments in the target code can mislead LLMs into producing irrelevant or incorrect tests. To address this, we introduce a code summarization step during preprocessing that replaces misleading comments with concise summaries, which are then embedded into the prompt. Last, since achieving maximum mutation score is inherently difficult, we further introduce an iterative test generation strategy to push LLMs toward killing more live mutants. Once the mutation score converges, we analyze both uncovered and live mutants, as well as specific mutation operators that remain particularly challenging for the LLM. Like many recent studies [72, 34, 60], our approach focuses on function-level testing. While this prior work also generates test cases for functions from complex datasets such as Defects4J [41], they handle cross-module dependencies explicitly by just providing them in the prompts. In contrast, our work does not aim to address dependency resolution at this stage, but instead focuses on making significant improvements in mutation scores for function-level testing. Nevertheless, as discussed below, one of our datasets includes solutions to challenging algorithmic problems with complex program structures, providing a more complex dataset for evaluating the effectiveness of MutGen in generating test cases.
@@
To evaluate MutGen, we use two datasets: HumanEval-Java, widely adopted in prior work [63, 23, 61], and a new LeetCode-Java corpus containing 100 randomly selected algorithmic problems with ground-truth solutions. The experimental results show that test cases generated by MutGen outperform both EvoSuite, a state-of-the-art search-based technique, and a vanilla LLM-based method in terms of mutation score. MutGen achieves mutation scores of 89.5% and 89.1% on HumanEval-Java and LeetCode-Java, respectively, across a total of 1,144 and 1,900 mutants.
@@
II Motivating Example and Informal Description
@@
When combined with the example code (including comments), LLMs can generate test cases that achieve high line and branch coverage. However, as demonstrated in prior work [28, 18, 9], high coverage does not necessarily imply strong fault detection capability, for example, when measured as mutation score. For instance, in our experiments, LLMs generate tests for the subject id_81 from HumanEval-Java with 100% line and branch coverage, yet the corresponding mutation score is only 4%.
Listing 2 shows a representative test function generated by Llama-3.3. In practice, multiple such test functions may be generated, often with a variety of data inputs aimed at maximizing code coverage. However, we observe that most of these inputs fail to cover corner cases such as February 29th on leap years, likely due to the absence of such examples in the provided comment block. The comments in the target code may also introduce additional issues. For example, the comment begins with “You have to write a function…”, which, when included as part of the prompt, may mislead the LLM into replicating the target code logic rather than generating a meaningful test function. Additionally, Llama-3.3 occasionally produces dates in incorrect formats such as dd-mm-yyyy, likely due to ambiguity or insufficient specification of the expected format in the comments provided within the program under test.
@@
which fails in both the original function and mutant. However, it frequently fails to generate:
@@
@Test
public void testDayBoundaryInApril() {
String date = ”04-01-2025”;
assertTrue(ValidDate.validDate(date));
@@
To address these issues, we propose MutGen, an approach that improves the fault detection capability of LLM-generated test cases via a two-stage process: preprocessing and generation, which are detailed in Section III-B and Sections III-C-III-E, respectively.
II-A Preprocessing Stage
@@
“This Java function, ‘validDate‘, validates a given date string in the format ”mm-dd-yyyy” and returns ‘true‘ if the date is valid according to specific rules regarding month and day ranges, and ‘false‘ otherwise. The input date string must be non-empty and follow the exact specified format. If any of these conditions are not met, the function will return ‘false‘.”
To address the second issue, we leverage mutation feedback to guide the LLM in generating test cases with higher fault detection capability. Before the generation stage, MutGen extracts mutation feedback from mutation reports produced by mutation testing tools. In this example, we use PITest [6] to generate a mutation report, as shown in Figure 3. From the report, MutGen retrieves key information, including the mutation location (line number), the mutant status (killed, survived, or uncovered), and the mutation operator applied to create the mutant. MutGen then integrates this feedback with the source code, forming a prompt (shown in Figure 4) that guides the LLM in generating more effective test cases during the generation stage. For example, the mutation feedback extracted from the report for the mutant shown in Listing 1 is as follows:
@@
II-B Generation Stage
Prompt Used for Generation # Unit Test Generator for Java using Junit4 ## Source Code Summary {{summary}} ## Source File {{source code without comments}} ## Mutation Feedback Analyze the live mutant to identify the defects; the generated test cases must look very different from the existing ones. 21. conditional boundary at line 24, uncovered: if (day<=1||day>30) return false; ## Output Format in YAML new_tests: - test_behavior: … test_name: … test_code: … new_imports: …
@@ II-B Generation Stage
To further enhance effectiveness, MutGen introduces an iterative generation strategy. In each iteration, additional test cases are generated to target the remaining live and uncovered mutants that were not killed by the previous test suite. After two iterations, for the running example, MutGen achieves a 100% mutation score. During this process, LLMs may occasionally generate test cases with method names that conflict with existing method names. For example, as shown in Listing 2, the LLM generates a test case for the boundary value 04-01-2025 in April. Later, it may generate another test case for 04-31-2025, also intended to test boundary conditions. However, both may be assigned the same method name, leading to naming conflicts. To resolve this, MutGen incorporates an additional guideline, specifically addressing method name conflicts, into the prompt used for fixing (shown in Figure 5).
In contrast, as demonstrated in Section IV, for this subject in HumanEval-Java, using the standard prompt introduced at the beginning of this section, the LLM can only produce a test suite achieving a 53% mutation score, which remains unchanged even after four iterations.
Prompt Used for Fixing ## Failed Test The following test failed in the test suite: {{ failed_test with error message }} You are processing execution failures. Please match only one of these guidelines and then try to correct the failed tests according to the given error messages. 1. Direct Invocation: • Always use assertion functions directly (e.g., assertEquals(value1, value2)) instead of prefixed calls like Assert.assertEquals(value1, value2). 2. Handling assertEquals Mismatches: • If an error follows the pattern: • test_file_name:line_number: expected:error_value but was:correct_value • And involves assertEquals, replace the expected value with the correct one. • Example: assertEquals(error_value, xxx); → assertEquals(correct_value, xxx); 3. Fixing assertTrue / assertFalse Errors: • If an error follows the pattern: • classname:line_number:null • And involves assertTrue or assertFalse, swap the function name while keeping all parameters unchanged. • Example: assertTrue(xxx); → assertFalse(xxx); 4. Handling ambiguous • If an error has the pattern: • reference to assertEquals is ambiguous • And involves assertEquals, add type casts to both parameters based on their respective types. • Example: assertEquals(expectedValue, actualValue); → assertEquals((long) expectedValue, (long) actualValue); 5. Method name conflicts • If an error follows the pattern: method is already defined in the class, • only append a number to the method name. • Example: public void test_null() → public void test_null1(); 6. Other errors • Try to fix the errors according to the error messages. ## Output Format in YAML The same applied to Figure 4.
@@
III Our Approach
Figure 6: MutGen Overview
In this section, we precisely describe and formalize our proposed approach, MutGen. As shown in Figure 6, MutGen takes a source program as input and returns test cases aimed at improving its fault-detection effectiveness by maximizing the test suite’s mutation score. We highlight all data flows in red in the figure. Steps 1, 2, and 5 are described in Sections III-B, III-C, and III-D, respectively. We do not include Section III-E as it reuses the components already shown in the figure. We begin with a high-level overview of MutGen, followed by detailed and formal descriptions of each individual component.
III-A Overview
Let $\mathcal{M}=\{m_{1},m_{2},\dots,m_{n}\}$ denote the set of mutants generated from the program under test. The mutation score $\mu(T)$ of a test suite $T$ is defined as:
@@ III-A Overview
The preprocessing stage prepares information to guide test case generation, including code summarization and mutation feedback, as detailed in Section III-B. In the generation stage, the LLM is prompted with mutation feedback, such as descriptions of mutation operators and diff-like representations of mutants, to generate test cases.
Let $\mathcal{R}=\{r_{1},r_{2},\dots,r_{k}\}$ denote the set of mutation feedback entries (i.e., killed, live, uncovered), where each $r_{i}$ is a tuple $\langle\text{id}_{i},m_{i},s_{i},\text{op}_{i}\rangle$, respectively representing the mutant identifier, the mutated statement, the mutant status, and the applied mutation operator. MutGen first generates an initial test suite $\mathcal{T}$:
@@
After executing $\mathcal{T}$, MutGen collects a set $\mathcal{F}=\{f_{1},...,f_{q}\}$ consisting of all test cases that encountered execution failures and corresponding error messages. For each $f_{i}$, MutGen constructs a second prompt $\text{Prompt}_{\text{fail}}$ (as illustrated in the example shown in Figure 5) and invokes the LLM to generate a set of fixed test cases $\mathcal{T}_{fix}$:
@@
This process will be detailed in Section III-D.
@@
Our assumption, as reported in the mutation testing literature [18, 11], is that by maximizing the mutation score, we also maximize fault detection effectiveness. In any case, achieving mutation coverage is known to be a more stringent criterion than line or branch coverage [28] and therefore enables the generation of more effective test suites. However, as revealed in recent work [32], finding an optimal prompt for LLMs remains a significant challenge, and we introduce an approximate and iterative solution, detailed in Section III-E, which we investigate empirically and whose limitations we analyze, for example, in terms of mutation operators that are difficult to kill.
III-B Preprocessing
@@
Code Summarization. As illustrated in Figure 6 and the example in Section II-A, the prompt instructs the LLM to summarize the purpose and the input format of the target code, resulting in a natural language description denoted as descdesc.
@@
As shown in the example in Section II-A, the corresponding mutation feedback entry is represented as:
@@
III-C Prompting with Mutation Feedback